OpenStudy (anonymous):

x-intercept of (2x+7)(x-1)

OpenStudy (noelgreco):

Find out where the numerator is equal to zero

OpenStudy (anonymous):

Assuming that you mean y = (2x+7)(x-1), set y to 0 and determine which values of x give that value.

OpenStudy (anonymous):

For example, the x-intercepts of the equation y = (x+a)(x-b) would be -a and b.

OpenStudy (anonymous):

hmm. I've tried to put in -7, 1 and 7, -1 and 7,1

OpenStudy (anonymous):

I'm really confused with this problem. Its really long and doesn't tell me if anything is wrong or right along the way. can i attach and you look?

OpenStudy (anonymous):

Since it's (2x+7), not (x+7), the x-intercept will not be -7.

OpenStudy (anonymous):

Think of it this way- if 2x + 7 = 0, what is the value of x?

OpenStudy (anonymous):

7?

OpenStudy (anonymous):

7/2?

OpenStudy (anonymous):

Not exactly, but close. Solve it like you would any other equation: \[2x+7 = 0\]\[2x = -7\]\[x = \frac{ -7 }{ 2 }\]

OpenStudy (anonymous):

Then, you do the same with the other part of the factored equation- what is x in x - 1=0?

OpenStudy (anonymous):

1

OpenStudy (anonymous):

So you have your x-intercepts then :)

OpenStudy (anonymous):

can you look at the attachment? it says the problem is wrong still. idk what is wrong though

OpenStudy (anonymous):

You're telling them that the intercepts are 7 and -1, not -7/2 and 1

OpenStudy (anonymous):

yea, i changed that part but it says its still wrong

OpenStudy (anonymous):

Maybe you have to put it in the form (x, 0)

OpenStudy (anonymous):

Or try putting it in decimal form.

OpenStudy (anonymous):

wont allow decimals.

OpenStudy (anonymous):

i have no idea

OpenStudy (anonymous):

I know it's right though, try graphing it: https://www.desmos.com/calculator

OpenStudy (anonymous):

well that part of the problem is right because the percentage i have went up

OpenStudy (anonymous):

but something else is wrong and idk what

OpenStudy (anonymous):

Looks all right to me. Must be the applet acting up.

OpenStudy (anonymous):

No wait.

OpenStudy (anonymous):

Maybe you're suppsoed to put the domain as [-inf, inf]?

OpenStudy (anonymous):

i dont think so. in the example it says to put it in as (-inf, inf)

OpenStudy (anonymous):

Well, either something's slightly off about the way you're imputting your answers, or the applet's messed up, because the answers are right.

OpenStudy (anonymous):

it might be the site because theres another problem that i'm almost positive is right

OpenStudy (anonymous):

this one